OpenStudy (anonymous):

multiply the following expression mentally 12 x 22= 12(20+2)=

OpenStudy (whpalmer4):

mentally write down your answer :-) 12*20 = 12*2 = those two added =

OpenStudy (whpalmer4):

you could also do this as 22(10+2) = 22*10 = (easy) 22*2 = (also easy) those two added = (same answer as the other way)